# Learning Successor Features with Distributed Hebbian Temporal Memory

### v2

@@ 1 Introduction
Modelling sequential data is one of the essential tasks in Artificial Intelligence as it has many applications, including decision-making and world models (Moerland et al., 2023), natural language processing (Min et al., 2021), conversational AI (Dwivedi et al., 2023), time-series analysis (Eraslan et al., 2019), and video and music generation (Ji et al., 2020). One of the classical approaches to modelling sequential data is forming a representation that stores and condenses the most relevant information about a sequence, and finding a general transformation rule of this information through the dimension of time (Lipton et al., 2015; Harshvardhan et al., 2020; Mathys et al., 2011). We refer to the class of algorithms that use this approach as Temporal Memory (TM) algorithms, as they essentially model the cognitive ability of complex living organisms to remember the experience and make future predictions based on this memory (Hochreiter and Schmidhuber, 1997; Friston et al., 2016, 2018; Parr and Friston, 2017).
@@
This paper addresses the problem of online sequence learning for planning and decision-making under uncertainty, which can be formalized as Reinforcement Learning (RL) for a Partially Observable Markov Decision Process (POMDP) (Poupart, 2005; Singh et al., 2021). Inferring the hidden state in a partially observable environment is, in effect, a sequence modelling problem as it requires processing a sequence of observations to get enough information about hidden states. One of the most efficient representations of the hidden states for discrete POMDP is the Successor Representation (SR) that disentangles hidden states and goals given by the reward function (Dayan, 1993; Samuel J. Gershman, 2018; Millidge and Buckley, 2022). The Successor Features framework is an extension of the SR into continuous POMDP, which employs the same idea of value function decomposition but, instead, for features of a hidden state (Barreto et al., 2017; Touati et al., 2022). Temporal Memory (TM) algorithms can be leveraged to make cumulative predictions about future states and their features, forming SR or SF. This work shows that the proposed algorithm, Distributed Hebbian Temporal Memory (DHTM), can effectively form SFs for navigation tasks in Gridworld and AnimalAI (Fountas et al., 2020) environments.
@@
The most prominent TM algorithms, like a Recurrent Neural Network (RNN) (Boden, 2002) or a Hidden Markov Model (HMM) (Eddy, 2004), use backpropagation to capture data relationships, known for their instability due to recurrent non-linear derivatives. They also require having complete sequences of data at hand during the training. Although the gradient vanishing problem can be partially circumvented in a way, Receptance Weighted Key Value (RWKV) (Peng et al., 2023) or Linear Recurrent Unit (LRU) (Orvieto et al., 2023) models do, the problem of online learning, i.e., learning on non-stationary datasets, is still a viable topic. In contrast to HMM, RNN models and their descendants also lack a probabilistic theory foundation, which is beneficial for modelling sequences captured from stochastic environments (Salaün et al., 2019; Zhao et al., 2020). There is little research on TM models that can be used in fully online adaptable systems interacting with partially observable stochastic environments with access only to one sequence data point at a time, a prevalent case in Reinforcement Learning (Jahromi et al., 2022).
@@
We propose a Distributed Hebbian Temporal Memory (DHTM) algorithm based on the factor graph formalism and multi-compartment neuron model. The resulting graphical structure of our model is similar to one of the Factorial-HMM (Ghahramani and Jordan, 1995) but with a factor graph forming online during training. An important feature of our model is that transition matrices for each factor are stored as different components (segments) of artificial neurons, which makes computations very efficient in the case of sparse transition matrices. Our TM forms sequence representations entirely online and employs only local Hebbian-like learning rules (Hebb, 2005; Churchland and Sejnowski, 1992; Lillicrap et al., 2020), circumventing gradient drawbacks and making the learning process much faster than gradient methods. This comes at the expense of hidden state generalization, as DHTM’s hidden state encodes each observation sequence uniquely. Different trajectories leading to the same position in partially observable Gridworld may result in different hidden states inferred by DHTM. It’s desirable if we want to store different experiences independently, like in human episodic memory (Zhang et al., 2021; W. Schulz and Robins, 2023). Moreover, we argue that generalization is unnecessary when the reward function is well-defined in observation or feature space, and it’s more efficient to use fast hash-like encodings.
@@
The DHTM model notoriously fits Successor Features in the Reinforcement Learning setup. The proposed TM is tested as a world model (Ha and Schmidhuber, 2018; Hafner et al., 2023) for an RL agent architecture, navigating in a Gridworld environment and a more challenging AnimalAI testbed (Crosby et al., 2020). Our algorithm outperforms a classic LSTM and a biologically-inspired HMM-like world model CSCG (George et al., 2021) in online Successor Feature formation task due to a combination of fast Hebbian-like learning and sparse hidden state coding. Another advantage of our algorithm is that it allows its implementation for neuromorphic processors, using only local learning rules.
Our contribution to this work is the following:
  • We propose a distributed memory model (DHTM) based on factor graph formalism and a multicompartment neural model.
  • Our model stores sparse factor functions in neural segments, which significantly lessens the number of trainable parameters, and speeds up learning.
@@
  • The DHTM learns fully online, employing only local Hebbian-like rules.
  • Tested as a world model for an RL agent architecture in Gridworld and AnimalAI, DHTM outperforms LSTM and CSCG in online Successor Features formation for navigation tasks.
@@ 2 Background
This section provides basic information about some concepts necessary to follow the paper.
@@ 2.1 Reinforcement Learning
This paper considers decision-making in a partially observable environment, usually formalized as a Partially Observable Decision Process (Poupart, 2005). A POMDP is defined as a tuple $\mathcal{M}=(S,A,P,R,O,D,\gamma)$, where $S$—state space, $A$—action space, $P(s,a,s`)=Pr(s`\ |\ s,a)$—transition function, $R(s)$–reward function, O—observation space, $D(a,s`,o)=Pr(o\ |\ a,s`)$—sensor model and $\gamma\in[1,0)$—discount factor, given a transition $s,a\to s`$, where $s\in S$, $a\in A$, $o\in O$. If $S,A,O$ are finite, $P,D$ can be viewed as real valued matrices, otherwise, they are conditional density functions. Here we consider deterministic rewards, which depend only on the current state, i.e. $R(s):S\to\mathbb{R}$.
@@ 2.2 Hidden Markov Model
Partially observable Markov process can be approximated by a Hidden Markov model (HMM) with hidden state space $H$ and observation space $O$. $O$ is the same as in $\mathcal{M}$, but $H$ generally is not equal $S$. Variables $H_{t}$ represent an unobservable (hidden) approximated state of the environment which evolves over time, and observable variables $O_{t}$ represent observations that depend on the same time step state $H_{t}$, and $h_{t},o_{t}$ are corresponding values of this random variables. For the sake of simplicity, we suppose that actions are fully observable and information about them is included into $H_{t}$ variables. For the process of length $T$ with state values $h_{1:T}=(h_{1},\ldots,h_{T})$ and $o_{1:T}=(o_{1},\ldots,o_{T})$, the Markov property yields the following factorization of the generative model:
@@
In case of discrete hidden state, a time-independent stochastic transition matrix can be learned with Baum–Welch algorithm (Baum et al., 1970), a variant of Expectation Maximization algorithm. To compute the statistics for the expectation step, it employs the forward-backward algorithm, which is a special case of sum-product algorithm (Kschischang et al., 2001).
2.3 Successor Representation
Successor Representations are such representations of hidden states from which we can linearly infer the state value given the reward function (Dayan, 1993). Here, we assume observation and state spaces are discrete.
$V(h_{t}=i)=\mathrm{E}[\sum^{\infty}_{l=0}\gamma^{l}R_{t+l+1}\ |\ h_{t}=i]=\\ =\sum^{\infty}_{l=0}\gamma^{l}\sum_{j}p(h_{t+l+1}=j\ |\ h_{t}=i)R_{j}=\\ =\sum_{j}\sum^{\infty}_{l=0}\gamma^{l}p(h_{t+l+1}=j\ |\ h_{t}=i)R_{j}=\sum_{j}M_{ij}R_{j},$ (2)
where $\gamma$ is a discount factor, vector $\mathrm{SR}(h=i)=\{M_{ij}\}_{j}$ is a Successor Representation of a state $i$, and $M_{ij}=\sum^{\infty}_{l=0}\gamma^{l}p(h_{t+l+1}=j\ |\ h_{t}=i)$. $R_{j}$ is a reward for observing the state $j$. That is, SR can be computed by a TM that is able to predict future states. TM algorithms effectively predict observations only for a finite time horizon $T$. Therefore, in order to learn SR, a technique similar to TD learning in standard RL may be employed:
$\delta_{ij}$ $=\sum^{T}_{l=0}\gamma^{l}p(h_{t+l+1}=j\ |\ h_{t}=i))+$ (3)
$+\gamma^{T+1}\sum_{k}M_{kj}p(h_{t+T+1}=k\ |\ h_{t}=i)-M_{ij},$
$M_{ij}$ $\leftarrow M_{ij}+\alpha\delta_{ij},$ (4)
where $\alpha\in(0,1)$ is a learning rate, $\delta_{ij}$—TD error for SR.
In partially observable environments, however, exact state values are not known, therefore we operate with state distributions or so-called belief states (Poupart, 2005), which are inferred from observations. In that case, state value and SR are functions of hidden state variable distribution (see details in Appendix B).
2.4 Sparse Distributed Representations
In our work, we design our model to operate with sparse distributed representations (SDRs) to reflect the spatiotemporal property of cortical network activity (Perin et al., 2011). In the discrete time case, SDR is a sparse binary vector in a high-dimensional space. To encode observed dense binary patterns to SDRs, we use a biologically plausible k-WTA (k-winners take all) neural network algorithm called spatial pooler with a Hebbian-like unsupervised learning method (see details in Appendix A).
@@ 3.1 Factor Graph Model
Distributed Hebbian Temporal Memory is based on the sum-product belief propagation algorithm in a factor graph (see Figure 1). Analogously to Factorial-HMM (Ghahramani and Jordan, 1997), we divide the hidden space $H$ into subspaces $H^{k}$. There are four sets of random variables (RV) in the model: $H^{i}_{t-1}$—latent variables representing hidden states from the previous time step (context), $H^{k}_{t}$—latent variables for the current time step, $\Phi^{k}_{t}$—feature variables, and $O^{lm}_{t}$—observable variables. Except for $O^{lm}_{t}$, all random variables have a categorical distribution. In contrast, $O^{lm}_{t}$, are Bernoulli variables because they represent pixels from a binary input image observation. RV state values are denoted as corresponding lowercase letters: $h^{i}_{t-1}$, $h^{k}_{t}$, $\varphi^{k}_{t}$, $o^{lm}_{t}$.
Each variable $\Phi^{k}_{t}$ is considered independent and has a separate graphical model for increased computational efficiency. However, hidden variables of the same time step are statistically interdependent in practice. We introduce their interdependence through a segment computation trick that goes beyond the standard sum-product algorithm (see Eq. 8).
The model also has three types of factors: $M^{i}_{t-1}$—messages from previous time steps, $F^{k}_{c}$—context factor (generalized transition matrix), $F^{k}_{e}$—emission factor. We assume that messages $M^{i}_{t-1}$ include posterior information from the time step $t-1$, therefore we don’t depict observable variables for previous time steps in Figure 1.
Figure 1: Partial factor graph for the DHTM. The input to the model is a sequence of binary images, each pixel is modelled as Bernoulli random variable $O^{lm}_{t}$, where $l$ and $m$ denote corresponding rows and cols of the image. The encoder block forms image categorical features $\Phi^{k}_{t}$ in an unsupervised manner. Each feature $\Phi$ has its own explaining hidden variable, which may depend on hidden variables of the other features and on itself from the previous time step. $F^{k}_{c}$ and $F^{k}_{e}$ are context and emission factors for the corresponding variables. Unary factors $M^{i}_{t-1}$ called messages represent accumulated information about previous time steps.
Further, we discuss only the upper block of the graph in Figure 1, which is DHTM itself. The lower block—an encoder—is described in the Appendix A. The only requirement for the encoder is that its output should be represented as states of categorical variables (features) for the current observation.
3.2 Neural Implementation
The main routine of the DHTM is to estimate distributions of currently hidden state variables given by the equation 5, the computational flow of which is schematically depicted in Figure 2:
@@ 3.2 Neural Implementation
Figure 2: Neuronal implementation of the DHTM. Random variables are represented by cell clusters (white circles), where each cell corresponds to a state and its spike frequency—to the probability of the state $p(h^{k}_{t})$. Cell’s dendritic segments $seg(k)$ correspond to context factor values $f_{l}$ for a particular combination of states (active presynaptic cells) $rec(l)$. Segments’ excitations $E_{l}$ are combined to determine cell’s spike frequency $p(h^{k}_{t})$. Segment’s synaptic weights reflect specificity of $rec(l)$ combination for the segment. Emission factors $F^{k}_{e}$ are fixed and represented by minicolumns inside a variable.
For computational purposes, we translate the problem to the neural network architecture with Hebbian-like learning (for biological interpretation of the model, see Appendix C). As can be seen from Figure 2, every RV can be viewed as a set of spiking neurons representing the RV’s states, that is, $p(h^{k}_{t})=p(c^{j}_{t}=1)$, where $j$—index of a neuron corresponding to the state $h^{k}_{t}$. Cell activity is binary $c^{j}_{t}\in\{0,1\}$ (spike/no-spike), and the probability might be interpreted as a spike rate. Factors $F^{k}_{c}$ and $M^{i}_{t-1}$ can be represented as vectors, where elements are factor values for all possible combinations of RV states included in the factor. Let’s denote elements of the vectors as $f_{l}$ and $m_{u}$ correspondingly, where $l$ corresponds to a particular combination of $k,h^{k}_{t},h^{i_{1}}_{t-1},\dots,h^{i_{n_{l}}}_{t-1}$ state values and $u$ indexes all neurons representing states of previous time step RVs.
Drawing inspiration from biological neural networks, we group a neuron’s connections into dendritic segments. A segment acts as an independent computational unit that detects a particular input pattern (a context state) defined by its own receptive field. In our model, a segment links together factor value $f_{l}$, the computational graph shown in Figure 2, and the excitation $E_{l}$ induced by the segment $l$ to the cell it is attached to. The segment is active, i.e., $s_{l}=1$ if all its presynaptic cells are active; otherwise, $s_{l}=0$. Computationally, a segment transmits its factor value $f_{l}$ to a cell it is attached to if the context matches the corresponding state combination.

### v3

@@ 1 Introduction
Modelling sequential data is one of the essential tasks in artificial intelligence as it has many applications, including decision-making and learning world models [1], natural language processing [2], conversational AI [3], time series analysis [4], and video and music generation [5]. One of the classical approaches to modelling sequential data is to form a representation that stores and condenses the most relevant information about a sequence, and to find a general transformation rule of this information through the dimension of time [6, 7, 8]. We refer to the class of algorithms that use this approach as Temporal Memory (TM) algorithms because they essentially model the cognitive ability of complex living organisms to remember the past experience and make future predictions based on that memory [9, 10, 11, 12].
@@
This paper addresses the problem of online sequence learning for planning and decision-making under uncertainty, which can be formalized as reinforcement learning (RL) for a partially observable Markov decision process (POMDP) [13, 14]. Inferring the hidden state in a partially observable environment is a sequence modelling problem, since it requires processing a sequence of observations to obtain enough information about the hidden state. One of the most efficient representations of hidden states for discrete POMDPs is the Successor Representation (SR), which disentangles hidden states and goals given by the reward function [15, 16, 17]. The Successor Features framework is an extension of SR to continuous POMDP, using the same idea of value function decomposition, but instead for features of a hidden state [18, 19]. Temporal memory (TM) algorithms can be used to make cumulative predictions about future states and their features, forming SR or SF. This work shows that the proposed algorithm, Distributed Hebbian Temporal Memory (DHTM), can dynamically form SFs for navigation tasks in Gridworld and AnimalAI [20] environments.
@@
The most prominent TM algorithms, such as a Recurrent Neural Network (RNN) [21] or LSTM [9], use backpropagation to capture data relationships, which is known to be unstable due to recurrent nonlinear derivatives. They also require complete data sequences to be available during training. Although the gradient vanishing problem can be partially circumvented in a way that Receptance Weighted Key Value (RWKV) [22] or Linear Recurrent Unit (LRU) [23] models do, the problem of online learning, i.e. learning on non-stationary data sets, is still a viable issue. Unlike graphical models such as HMM [24], RNN models and their descendants also lack a probabilistic theoretical foundation, which is advantageous for modelling sequences captured from stochastic environments [25, 26]. There is little research on TM models that can be used in fully online adaptive systems interacting with partially observable stochastic environments with access to only one sequence data point at a time, a common case in reinforcement learning [27].
@@
We propose a Distributed Hebbian Temporal Memory (DHTM) algorithm based on factor graph formalism and a multi-compartment neuron model inspired by the HTM neuron [28]. The resulting graphical structure of our model is similar to that of the factorial HMM [29]. An important feature of our model is that the transition matrices for each factor are stored as different components (segments) of artificial neurons, which makes computation very efficient in the case of sparse transition matrices. Our TM forms sequence representations entirely online and uses only local Hebbian-like learning rules [30, 31, 32], which avoid the drawbacks of gradient methods and make the learning process much faster than gradient methods. This comes at the expense of hidden state generalization, since DHTM’s hidden state uniquely encodes each observation sequence. Different trajectories leading to the same position in the partially observable Gridworld can result in different hidden states inferred by DHTM. In essence, this makes DHTM a model of episodic memory [33], or an adaptive trajectory buffer, with built-in machinery for making multistep future predictions. In this light, we also show that an agent forming SF with DHTM can be interpreted as a kind of episodic control [34, 35, 36, 37].
@@
The proposed TM is tested as a world model for an RL agent architecture navigating in a Gridworld environment and a more challenging AnimalAI testbed [38]. The results demonstrate that our algorithm outperforms classical LSTM and a biologically inspired HMM-like world model CSCG [39] in changing environments and is in order of magnitude more sample efficient than Dreamer V3 [40]. Another advantage of our algorithm is that it can be implemented on neuromorphic processors using only local learning rules.
Our contribution to this work includes the following key elements:
  • A novel episodic memory model, called DHTM, which is based on a factor graph formalism and features an efficient, bio-inspired implementation.
@@
  • We demonstrate that the DHTM can operate fully online, using only local Hebbian-like learning rules.
  • We show that DHTM learns at a speed comparable to model-based table episodic control while supporting distributed representations. This capability significantly broadens its applicability.
@@ 2 Background
This section provides basic information about some of the concepts necessary to follow the paper.
@@ 2.1 Reinforcement Learning
This paper considers decision-making in a partially observable environment, usually formalized as a partially observable decision process [13]. A POMDP is defined as a tuple $\mathcal{M}=(S,A,P,R,O,D,\gamma)$, where $S$—state space, $A$—action space, $P(s,a,s`)=Pr(s`\ |\ s,a)$—transition function, $R(s)$–reward function, O—observation space, $D(a,s`,o)=Pr(o\ |\ a,s`)$—sensor model and $\gamma\in[1,0)$—discount factor, given a transition $s,a\to s`$, where $s\in S$, $a\in A$, $o\in O$. If $S,A,O$ are finite, $P,D$ can be viewed as real-valued matrices, otherwise, they are conditional density functions. Here we consider deterministic rewards that depend only on the current state, i.e. $R(s):S\to\mathbb{R}$.
@@
One of the drawbacks of value-based methods is that they require learning a new Q-function when the reward function changes, even if the environment remains the same. To get around this, the dynamics of the environment and the reward function can be decoupled, which leads us to the successor representation framework.
2.2 Successor Representation
Successor representations are such representations of hidden states from which we can linearly infer the state value given the reward function [15]. We assume that the observation and state spaces are discrete.
$V(h_{t}=i)=\mathrm{E}[\sum^{\infty}_{l=0}\gamma^{l}R_{t+l+1}\ |\ h_{t}=i]=\sum_{j}\mathrm{SR}_{ij}R_{j},$ (1)
where $\gamma$ is a discount factor, $i$-th row in the matrix $\mathrm{SR}$ is the Successor Representation of a state $i$, and $\mathrm{SR}_{ij}=\sum^{\infty}_{l=0}\gamma^{l}p(h_{t+l+1}=j\ |\ h_{t}=i)$. $R_{j}$ is a reward for observing the state $j$.
An extension of SR for large or continuous state spaces that we use is Successor Features [18], which is based on a similar idea of value function decomposition, but in a feature space instead (see details in Appendix A).
It is easy to show that, similarly to the Q-function, both SR and SF can be obtained using dynamic programming. However, this limits its use to relatively stable environments, since changes in the transition function can’t be taken into account immediately.
The next step to improve an agent’s adaptability is to use a world model to approximate the environment’s transition function and use it directly to form SR according to its definition. And one of the most basic such world models is the Hidden Markov Model.
@@ 2.3 Hidden Markov Model
A partially observable Markov process can be approximated by a Hidden Markov Model (HMM) with hidden state space $H$ and observation space $O$. $O$ is the same as in $\mathcal{M}$, but $H$ is generally not equal to $S$. Variables $H_{t}$ represent an unobservable (hidden) approximate state of the environment that evolves over time, and observable variables $O_{t}$ represent observations that depend on the same time-step state $H_{t}$, and $h_{t},o_{t}$ are corresponding values of these random variables. For simplicity, we assume that actions are fully observable and that information about them is contained in $H_{t}$ variables. For the process of length $T$ with state values $h_{1:T}=(h_{1},\ldots,h_{T})$ and $o_{1:T}=(o_{1},\ldots,o_{T})$, the Markov property yields the following factorization of the generative model:
@@
In the case of a discrete hidden state variable, a time-independent stochastic transition matrix can be learned with the Baum-Welch algorithm [41], a variant of the expectation maximization (EM) algorithm. To compute the statistics for the expectation step, it uses the forward-backward algorithm, which is a special case of the sum-product algorithm [42].
Since the transition matrix grows quadratically with the size of the hidden state space, it may become infeasible to train the HMM in large state spaces. Since it is possible that not all hidden states are dependent on each other, the Factorial HMM can mitigate this problem by partitioning the hidden state spaces into smaller independent subspaces, providing an efficient way to handle distributed representations.
We base DHTM on ideas from the Baum-Welch algorithm and the Factorial HMM, while adapting them for online learning.
2.4 Episodic Control
Episodic Control (EC) is a general computational approach to storing and reusing an agent’s experience, inspired by the notion of episodic memory [33] from psychology.
[…]
In this paper, we refer to EC to better explain the principle of the proposed algorithm for solving RL tasks.
@@ 3.1 Factor Graph Model
Distributed Hebbian Temporal Memory is based on the sum-product belief propagation algorithm in a factor graph (see Figure 1A). Analogous to Factorial HMM [44], we divide the hidden space $H$ into subspaces $H^{k}$. There are four sets of random variables (RV) in the model: $H^{i}_{t-1}$—latent variables representing hidden states from the previous time step (context), $H^{k}_{t}$—latent variables for the current time step, and $\Phi^{k}_{t}$–feature variables. All random variables have categorical distributions. RV state values are denoted by corresponding lowercase letters: $h^{i}_{t-1}$, $h^{k}_{t}$, $\varphi^{k}_{t}$, $o^{lm}_{t}$.
Each variable $\Phi^{k}_{t}$ is considered independent and has a separate graphical model for increased computational efficiency. However, in practice, hidden variables of the same time step are statistically interdependent. We introduce their interdependence through a segment computation trick that goes beyond the standard sum-product algorithm (see equation 6).
The model also has three types of factors: $M^{i}_{t-1}$—messages from previous time steps, $F^{k}_{c}$—context factor (generalized transition matrix), $F^{k}_{e}$–emission factor. We assume that messages $M^{i}_{t-1}$ contain posterior information from time step $t-1$, so we don’t plot observable feature variables for previous time steps in Figure 1A.
Figure 1: A. Example of a factor graph for the DHTM. The input to the model is a sequence of observations $o_{t}$. The encoder block forms categorical features $\Phi^{k}_{t}$. Each feature $\Phi$ has its own explaining hidden variable, which may depend on hidden variables of the other features and on itself from the previous time step. $F^{k}_{c}$ and $F^{k}_{e}$ are context and emission factors for the corresponding variables. Unary factors $M^{i}_{t-1}$ called messages represent accumulated information about previous time steps. B. Neural implementation. Random variables are represented by cell clusters (white circles), where each cell corresponds to a state and its spike frequency to the probability of the state $p(h^{k}_{t})$. The dendritic segments of the cell correspond to the context factor values $f_{l}$ for a given combination of states (active presynaptic cells). Segments’ excitations $E_{l}$ are combined to determine the cell’s spike frequency $p(h^{k}_{t})$. Segment synaptic weights reflect the specificity of the combination of presynaptic cells for the segment. Emission factors $F^{k}_{e}$ are fixed and represented by minicolumns within a variable.
We will discuss only the upper block of the graph in Figure 1A, which is DHTM itself. The lower block—an encoder—is described in Appendix C. The only requirement for the encoder is that its output should be represented as states of categorical variables (features) for the current observation.
The main routine of the DHTM is to estimate distributions of hidden state variables given by the equation 3, the computational flow of which is schematically shown in Figure 1B:
@@
3.2 Neural Implementation
For computational purposes, we translate the factor graph to the Hebbian learning neural network architecture (for a biological interpretation of the model, see Appendix B). As shown in Figure 1B, each RV can be viewed as a set of spiking neurons representing the states of the RV, i.e. $p(h^{k}_{t})=p(c^{j}_{t}=1)$, where $j$–is the index of a neuron corresponding to the state $h^{k}_{t}$. Cell activity is binary $c^{j}_{t}\in\{0,1\}$ (spike/no spike), and the probability can be interpreted as a spike rate. Factors $F^{k}_{c}$ and $M^{i}_{t-1}$ can be represented as vectors, where elements are factor values for all possible combinations of RV states included in the factor. Let’s denote the elements of the vectors as $f_{l}$ and $m_{u}$ respectively, where $l$ corresponds to a particular combination of $k$ state values $h^{k}_{t},h^{i_{1}}_{t-1},\dots,h^{i_{n_{l}}}_{t-1}$, and $u$ indexes all neurons representing states of RVs from previous time steps.
Inspired by biological neural networks, we group the connections of a neuron into dendritic segments. A segment acts as an independent computational unit that detects a particular input pattern (a contextual state) defined by its receptive field. In our model, a segment connects the factor value $f_{l}$ and the excitation $E_{l}$ induced by the segment $l$ to the cell to which it is connected. The segment is active, i.e. $s_{l}=1$ if all its presynaptic cells are active, otherwise $s_{l}=0$. Computationally, a segment transmits its factor value $f_{l}$ to the cell to which it is attached if the context matches the corresponding state combination.